# Retrieval Enhanced Zero-Shot Video Captioning

### v1

@@ 4.5. Qualitative Results
5. Conclusion
In this paper, we make the first attempt to explore the combination of a video-text retrieval model, an image-text matching model, and a large language model under the zero-shot and fast adaptation setting for video captioning. Specifically, we use learnable tokens as an information medium connecting these frozen models. With the help of our carefully designed loss functions, these tokens are able to convey video information to the LLM in a way that serves as suitable soft prompts. The pseudo-target-based inference-time token update ensures fast adaptation. Our proposed framework is conceptually simple and easy to implement. The extensive experiments on three benchmarks demonstrate its favorable zero-shot ability. We will make the code publicly available and hope our work can inspire future research in other areas.

### v2

@@ 4.5 Qualitative Results
Figure 8: Several video captioning qualitative examples using different methods.
Figure 9: Several image captioning qualitative examples using different methods.
5 Conclusion
In this paper, we make the first attempt to explore the combination of a video-text retrieval model, an image-text matching model, and a large language model under the zero-shot and fast adaptation setting for video captioning. Specifically, we use learnable tokens as an information medium connecting these frozen models. With the help of our carefully designed loss functions, these tokens are able to convey video information to the LLM in a way that serves as suitable soft prompts. The soft target-based inference-time token update ensures fast adaptation. Our proposed framework is conceptually simple and easy to implement. The extensive experiments on three benchmarks demonstrate its favorable zero-shot ability. We will make the code publicly available and hope our work can inspire future research in other areas.